NI TCS last activity

Go down

NI TCS last activity

Post  Alessio on Fri Feb 19, 2010 4:23 pm

(Alessio, Viviana, F. Frasconi, F. Gherardini, P. Ruggi)

Today, we finished to fix the mirror B holder.
Then we re-connected the cables of the stepping motors to mirror
B and checked that all the motors are working and the mirror is moving.
Finally we also checked the end-stops of the motors. The screen used
to position the cross-hair on the ITM is still in place.
The re-installation of the baffles will be done by Franco Frasconi
next week, right after all the other operations on the payload are
completed.
avatar
Alessio

Posts : 127
Join date : 2008-11-07
Age : 46
Location : Rome

View user profile

Back to top Go down

Back to top

- Similar topics

 
Permissions in this forum:
You cannot reply to topics in this forum